# Finite Neural Networks as Mixtures of Gaussian Processes: From Provable Error Bounds to Prior Selection

### v1

@@ A.2 Proofs Section 5
where $\mathcal{N}(z\mid-c_{i},1)$ denotes the pdf of a Gaussian distribution with mean $-c_{i}$ and variance $1$, and ${\mathbf{1}}_{[l,u]}(z)$ is the indicator function which equals $1$ if $z\in[l,u]$ and 00 otherwise. Note that, $\frac{1}{\Phi_{i}}{\mathbf{1}}_{[l_{i}-c_{i},u_{i}-c_{i}]}(z)\mathcal{N}(z\mid-c_{i},1)$ is in fact the pdf of a random variable with a Gaussian distribution $\mathcal{N}(-c_{i},1)$ conditioned on $[l_{i}-c_{i},u_{i}-c_{i}]$. Hence, the above conditional expectation is equivalent to the second moment of a truncated Gaussian random variable on $[l_{i}-c_{i},u_{i}-c_{i}]$, for which closed-form expressions exist (Wilhelm et al., 2012). Following Section 3 of Wilhelm et al. (2012) we can write:
[…]
Thus, we finally obtain:
$\sum_{i=1}^{N}\Phi_{i}{\mathbb{E}}_{{{z}}\sim\mathcal{N}(0,1)}[\|{{z}}-c_{i}\|^{2}\mid{{z}}\in[l_{i},u_{i}]]=\sum_{i=1}^{N}\nu_{i}+(\mu_{i}-c_{i})^{2}.$
Corollary 10 of Proposition 9: Wasserstein Distance from Signatures of Multivariate Gaussians
Proof  According to Proposition 8, we have:
@@
${\mathbb{W}}_{2}^{2}({\Delta}_{\mathcal{C}}\#\mathcal{N}({{\bar{0}}},\Sigma),\mathcal{N}({{\bar{0}}},\Sigma))$ $=\sum_{i=1}^{N}\mathbb{P}_{{{z}}\sim\mathcal{N}({{\bar{0}}},\Sigma)}[{{z}}\in{\mathcal{R}}_{i}]{\mathbb{E}}_{{{z}}\sim\mathcal{N}({{\bar{0}}},\Sigma)}[\|{{z}}-{{c}}_{i}\|^{2}\mid{{z}}\in{\mathcal{R}}_{i}].$
For all $i\in\{1,\ldots,N\}$, we can write
$\mathbb{P}_{{{z}}\sim\mathcal{N}({{\bar{0}}},\Sigma)}[{{z}}\in{\mathcal{R}}_{i}]{\mathbb{E}}_{{{z}}\sim\mathcal{N}({{\bar{0}}},\Sigma)}[\|{{z}}-{{c}}_{i}\|^{2}\mid{{z}}\in{\mathcal{R}}_{i}]$
[…]
By applying substitution ${{q}}={{T}}{{z}}$ to the integral terms, we obtain:
[…]
Here, since ${\mathcal{C}}$ is a grid in the transformed space induced by ${{T}}$, $\{{{T}}{{z}}\mid{{z}}\in{\mathcal{R}}_{i}\}$ is an axis-aligned hyper-rectangle that can be defined by vectors ${{l}}_{i},{{u}}_{i}\in{\mathbb{R}}^{n}$, i.e., $\{{{T}}{{z}}\mid{{z}}\in{\mathcal{R}}_{i}\}=[{{l}}_{i},{{u}}_{i}]$. Furthermore, because matrix ${{V}}$ is orthogonal, we have:
$\|{{S}}{{q}}\|^{2}={{q}}^{T}(\mathrm{diag}({{\lambda}})^{-\frac{1}{2}}{{V}}^{T}{{V}}\mathrm{diag}({{\lambda}})^{-\frac{1}{2}})^{-1}{{q}}={{q}}^{T}\mathrm{diag}({{\lambda}}){{q}}=\sum_{l=1}^{n}{{{\lambda}}^{(l)}}{{{q}}^{(l)}}^{2}.$
Hence, we can write
[…]
where $\bar{{{c}}}_{i}={{T}}{{c}}_{i}$. Using that $\mathcal{N}({{q}}\mid{{\bar{0}}},{{I}})=\prod_{l=1}^{n}\mathcal{N}({{{q}}^{(l)}}\mid 0,1)$, we can further simplify the above expression as follows
[…]
$\qquad=\sum_{l=1}^{n}{{{\lambda}}^{(l)}}\Big{(}\prod_{\begin{subarray}{c}j=1\end{subarray}}^{n}\mathbb{P}_{{{{q}}^{(j)}}\sim\mathcal{N}(0,1)}[{{{q}}^{(j)}}\in[{{{l}}_{i}^{{(j)}}},{{{u}}_{i}^{{(j)}}}]]\Big{)}{\mathbb{E}}_{{{{q}}^{(l)}}\sim\mathcal{N}(0,1)}[({{{q}}^{(l)}}-{{{c}}_{i}^{{(l)}}})^{2}\mid{{{q}}^{(l)}}\in[{{{l}}_{i}^{{(l)}}},{{{u}}_{i}^{{(l)}}}]]$
$\qquad=\Big{(}\prod_{\begin{subarray}{c}j=1\end{subarray}}^{n}\underbrace{\mathbb{P}_{{{{q}}^{(j)}}\sim\mathcal{N}(0,1)}[{{{q}}^{(j)}}\in[{{{l}}_{i}^{{(j)}}},{{{u}}_{i}^{{(j)}}}]]}_{\eqqcolon{\mathbb{P}_{i}^{{(j)}}}}\Big{)}\sum_{l=1}^{n}{{{\lambda}}^{(l)}}\underbrace{{\mathbb{E}}_{{{{q}}^{(l)}}\sim\mathcal{N}(0,1)}[({{{q}}^{(l)}}-{{{c}}_{i}^{{(l)}}})^{2}\mid{{{q}}^{(l)}}\in[{{{l}}_{i}^{{(l)}}},{{{u}}_{i}^{{(l)}}}]]}_{\eqqcolon{{\mathbb{E}}_{i}^{{(j)}}}}$
Because of Eqn. (10), without loss of generality, we can set the indexing of the signature locations, and hence the related regions, such that for all $i\in N$:
$\bar{{{c}}}_{i}=\bar{{{c}}}_{i_{1}\cdot i_{2}\cdot\ldots\cdot i_{n}}=({\bar{{{c}}}_{i_{1}}^{{(1)}}},{\bar{{{c}}}_{i_{2}}^{{(2)}}},\ldots,{\bar{{{c}}}_{i_{n}}^{{(n)}}})^{T},$
where $i_{l}\in\{1,\ldots,|{\mathcal{C}}^{l}|\}$, and ${\bar{{{c}}}_{i_{l}}^{{(l)}}}\in{\mathcal{C}}^{l}$, $\forall l\in\{1,\ldots,n\}$. Using straightforward algebra, we can now write
$\sum_{i=1}^{N}\Big{(}\prod_{j=1}^{n}{\mathbb{P}_{i}^{{(j)}}}\Big{)}\Big{(}\sum_{l=1}^{n}{{\mathbb{E}}_{i}^{{(l)}}}\Big{)}$ $=\sum_{i_{1}=1}^{|{\mathcal{C}}^{1}|}\sum_{i_{2}=1}^{|{\mathcal{C}}^{2}|}\ldots\sum_{i_{n}=1}^{|{\mathcal{C}}^{n}|}\Big{(}\prod_{j=1}^{n}{\mathbb{P}_{i_{1}\cdot i_{2}\ldots\cdot i_{n}}^{{(j)}}}\Big{)}\Big{(}\sum_{l=1}^{n}{{\mathbb{E}}_{i_{1}\cdot i_{2}\ldots\cdot i_{n}}^{{(l)}}}\Big{)}$
$=\sum_{i_{1}=1}^{|{\mathcal{C}}^{1}|}{\mathbb{P}_{i_{1}\cdot i_{2}\ldots\cdot i_{n}}^{{(1)}}}\sum_{i_{2}=1}^{|{\mathcal{C}}^{2}|}{\mathbb{P}_{i_{1}\cdot i_{2}\ldots\cdot i_{n}}^{{(2)}}}\ldots\sum_{i_{n}=1}^{|{\mathcal{C}}^{n}|}{\mathbb{P}_{i_{1}\cdot i_{2}\ldots\cdot i_{n}}^{{(n)}}}\sum_{l=1}^{n}{{\mathbb{E}}_{i_{1}\cdot i_{2}\ldots\cdot i_{n}}^{{(l)}}}$
$=\sum_{l=1}^{n}\Big{(}\prod_{\begin{subarray}{c}k=1\\ k\neq l\end{subarray}}^{n}\sum_{i_{k}=1}^{|{\mathcal{C}}^{k}|}{\mathbb{P}_{i_{1}\cdot i_{2}\ldots\cdot i_{n}}^{{(k)}}}\Big{)}\Big{(}\sum_{i_{l}=1}^{|{\mathcal{C}}^{l}|}{\mathbb{P}_{i_{1}\cdot i_{2}\ldots\cdot i_{n}}^{{(l)}}}{{\mathbb{E}}_{i_{1}\cdot i_{2}\ldots\cdot i_{n}}^{{(l)}}}\Big{)}$
Note that, as a consequence of our indexing, $\cup_{i_{k}=1}^{|{\mathcal{C}}^{k}|}[{{{l}}_{\ldots\cdot i_{k}\cdot\ldots}^{{(k)}}},{{{u}}_{\ldots\cdot i_{k}\cdot\ldots}^{{(k)}}}]={\mathbb{R}}$, and hence
$\sum_{i_{k}=1}^{|{\mathcal{C}}^{k}|}{\mathbb{P}_{\ldots\cdot i_{k}\cdot\ldots}^{{(k)}}}=\sum_{i_{k}=1}^{|{\mathcal{C}}^{k}|}\mathbb{P}_{{{{q}}^{(k)}}\sim\mathcal{N}(0,1)}[{{{q}}^{(k)}}\in[{{{l}}_{\ldots\cdot i_{k}\cdot\ldots}^{{(k)}}},{{{u}}_{\ldots\cdot i_{k}\cdot\ldots}^{{(k)}}}]]=1.$
Furthermore for $l\in\{1,\ldots,n\}$,
$\sum_{i_{l}=1}^{|{\mathcal{C}}^{l}|}{\mathbb{P}_{\ldots\cdot i_{l}\cdot\ldots}^{{(l)}}}{{\mathbb{E}}_{\ldots\cdot i_{k}\cdot\ldots}^{{(l)}}}$
$\qquad={{{\lambda}}^{(l)}}\sum_{i_{l}=1}^{|{\mathcal{C}}^{k}|}\mathbb{P}_{{{{q}}^{(l)}}\sim\mathcal{N}(0,1)}[{{{q}}^{(l)}}\in[{{{l}}_{\ldots\cdot i_{l}\cdot\ldots}^{{(l)}}},{{{u}}_{\ldots\cdot i_{l}\cdot\ldots}^{{(l)}}}]]{\mathbb{E}}_{{{{q}}^{(l)}}\sim\mathcal{N}(0,1)}[({{{q}}^{(l)}}-{\bar{{{c}}}_{\ldots\cdot i_{l}\cdot\ldots}^{{(l)}}})^{2}\mid{{{q}}^{(l)}}\in[{{{l}}_{\ldots\cdot i_{l}\cdot\ldots}^{{(l)}}},{{{u}}_{\ldots\cdot i_{l}\cdot\ldots}^{{(l)}}}]]$
$\qquad={{{\lambda}}^{(l)}}{\mathbb{W}}_{2}^{2}({\Delta}_{{\bar{{\mathcal{C}}}^{{(l)}}}}\#\mathcal{N}(0,1),\mathcal{N}(0,1)).$
Combining the above results, we obtain:

### v2

@@ B.1 Proofs of the Results in Section 5
B.2 Proofs of the Results in Section 6
Proofs of the ${\mathbb{W}}_{2}$ Distance Bounds in Table 13
The bound for the signature and compression operations directly follows from the triangle inequality. The bound for the affine transformation is a direct application of Lemma 32, as is the bound on ${\mathbb{W}}({\sigma}\#p,{\sigma}\#q)$.
@@ B.2 Proofs of the Results in Section 6
To derive the bound on ${\mathbb{W}}^{2}_{2}({\sigma}\#p,{\sigma}\#{\Delta}_{\mathcal{C}}\#p)$, we invoke Proposition 9, which yields
Since the activation function $\sigma$} is ${\mathcal{L}}_{{\sigma}\mid{\mathcal{R}}_{i}}$-Lipschitz on region ${\mathcal{R}}_{i}$, we have $\|{\sigma}({{z}})-{\sigma}({{c}}_{i})\|\leq{\mathcal{L}}_{{\sigma}\mid{\mathcal{R}}_{i}}\|{{z}}-{{c}}_{i}\|$ for all ${{z}},{{c}}_{i}\in{\mathcal{R}}_{i}$. Substituting this bound gives
Proof of Theorem 17: ${\mathbb{W}}_{2}$-Error of the GMM Approximation of a SNN at a Single Input Point
We prove Theorem 17 via induction and generalize the result to the $\rho$-Wasserstein distance with $\rho\in\{1,2\}$. To do so, we first derive the base case of the induction. For $k\in\{1,\ldots,K\}$, we have that:
${\mathbb{W}}_{\rho}\left(p_{nn({{x}}),k+1},q_{nn({{x}}),k+1}\right)$
(By Eqns (4) & (19))
$={\mathbb{W}}_{\rho}\left({\mathbb{E}}_{{{{z}}}_{k}\sim p_{nn({{x}}),k}}\left[L^{{{w}}_{k}}({\sigma}({{{z}}}_{k}))\#p_{{{w}}_{k}}\right],{\mathbb{E}}_{\tilde{{{{z}}}}_{k}\sim{G}_{{M_{k}}}(q_{nn({{x}}),k})}\left[L^{{{w}}_{k}}({\sigma}({\Delta}_{{\mathcal{C}}_{k}}(\tilde{{{{z}}}}_{k})))\#p_{{{w}}_{k}}\right]\right)$
(By the bound for the affine operation in Table 13))
$\leq{\mathbb{E}}_{{{w}}_{k}\sim p_{{{w}}_{k}}}\left[{\mathcal{L}}_{L^{{{w}}_{k}}}^{\rho}\right]^{1/\rho}{\mathbb{W}}_{\rho}\left({\sigma}\#p_{nn({{x}}),k},{\sigma}\#{\Delta}_{{\mathcal{C}}_{k}}\#{G}_{{M_{k}}}(q_{nn({{x}}),k})\right)$
(By the triangle inequality)
$\leq{\mathbb{E}}_{{{w}}_{k}\sim p_{{{w}}_{k}}}\left[{\mathcal{L}}_{L^{{{w}}_{k}}}^{\rho}\right]^{1/\rho}\left[{\mathbb{W}}_{\rho}\left({\sigma}\#p_{nn({{x}}),k},{\sigma}\#q_{nn({{x}}),k}\right)+{\mathbb{W}}_{\rho}\left({\sigma}\#q_{nn({{x}}),k},{\sigma}\#{G}_{{M_{k}}}(q_{nn({{x}}),k})\right)\right$.
$\qquad\left.+{\mathbb{W}}_{\rho}\left({\sigma}\#{G}_{{M_{k}}}(q_{nn({{x}}),k}),{\sigma}\#{\Delta}_{{\mathcal{C}}_{k}}\#{G}_{{M_{k}}}(q_{nn({{x}}),k})\right)\right]$
(By Lemma 32, i.e., applying the bound for the activation operation in Table 13)
$\leq{\mathbb{E}}_{{{w}}_{k}\sim p_{{{w}}_{k}}}\left[{\mathcal{L}}_{L^{{{w}}_{k}}}^{\rho}\right]^{1/\rho}{\mathcal{L}}_{\sigma}\left[{\mathbb{W}}_{\rho}\left(p_{nn({{x}}),k},q_{nn({{x}}),k}\right)+{\mathbb{W}}_{\rho}\left(q_{nn({{x}}),k},{G}_{{M_{k}}}(q_{nn({{x}}),k})\right)\right$.
$\qquad\left.+{\mathbb{W}}_{\rho}\left({G}_{{M_{k}}}(q_{nn({{x}}),k}),{\Delta}_{{\mathcal{C}}_{k}}\#{G}_{{M_{k}}}(q_{nn({{x}}),k})\right)\right]$
(By the definition of the ${\mathbb{MW}}_{2}$-distance, and because ${\mathbb{W}}_{1}\leq{\mathbb{W}}_{2}$)
$\leq{\mathbb{E}}_{{{w}}_{k}\sim p_{{{w}}_{k}}}\left[{\mathcal{L}}_{L^{{{w}}_{k}}}^{\rho}\right]^{1/\rho}{\mathcal{L}}_{\sigma}\left[{\mathbb{W}}_{\rho}\left(p_{nn({{x}}),k},q_{nn({{x}}),k}\right)+{\mathbb{MW}}_{2}\left(q_{nn({{x}}),k},{G}_{{M_{k}}}(q_{nn({{x}}),k})\right)\right$.
$\qquad\left.+{\mathbb{W}}_{\rho}\left({G}_{{M_{k}}}(q_{nn({{x}}),k}),{\Delta}_{{\mathcal{C}}_{k}}\#{G}_{{M_{k}}}(q_{nn({{x}}),k})\right)\right]$
(By Eqn (14) and because ${\mathbb{W}}_{1}\leq{\mathbb{W}}_{2}$)
$\leq{\mathbb{E}}_{{{w}}_{k}\sim p_{{{w}}_{k}}}\left[{\mathcal{L}}_{L^{{{w}}_{k}}}^{\rho}\right]^{1/\rho}{\mathcal{L}}_{\sigma}\left[{\mathbb{W}}_{\rho}\left(p_{nn({{x}}),k},q_{nn({{x}}),k}\right)+{\mathbb{MW}}_{2}\left(q_{nn({{x}}),k},{G}_{{M_{k}}}(q_{nn({{x}}),k})\right)\right$.
$\qquad\left.+\hat{\mathbb{W}}_{2,\bm{{\Delta}}_{\bm{{\mathcal{C}}}_{k}}}\left({G}_{M_{k}}(q_{nn({{x}}),k})\right)\right]$
Hence, if ${\mathbb{W}}_{\rho}\left(p_{nn({{x}}),k},q_{nn({{x}}),k}\right)\leq\hat{\mathbb{W}}_{\rho,k}$, then
${\mathbb{W}}_{\rho}\left(p_{nn({{x}}),k+1},q_{nn({{x}}),k+1}\right)$ $\leq{\mathbb{E}}_{{{w}}_{k}\sim p_{{{w}}_{k}}}\left[{\mathcal{L}}_{L^{{{w}}_{k}}}^{\rho}\right]^{1/\rho}{\mathcal{L}}_{\sigma}\left[\hat{{\mathbb{W}}}_{\rho,k}+{\mathbb{MW}}_{2}\left(q_{nn({{x}}),k},{G}_{{M_{k}}}(q_{nn({{x}}),k})\right)\right$.
$\qquad\left.+\hat{\mathbb{W}}_{2,\bm{{\Delta}}_{\bm{{\mathcal{C}}}_{k}}}\left({G}_{M_{k}}(q_{nn({{x}}),k})\right)\right]=\hat{{\mathbb{W}}}_{\rho,k+1}$
${\mathbb{W}}_{\rho}\left(p_{nn({{x}})},q_{nn({{x}})}\right)\leq\hat{{\mathbb{W}}}_{\rho,K+1}.$
Proof of Corollary 20: ${\mathbb{W}}_{2}$-Error of the GMM Approximation of a SNN at a Set of Input Points
The corollary follows directly from Theorem 17 by replacing $L^{{{w}}_{k}}$ with the stacked affine operation
$\bm{L}^{{{w}}_{k}}\left(({{z}}_{1}^{T},\ldots,{{z}}_{D}^{T})^{T}\right)=\left(L^{{{w}}_{k}}({{z}}_{1})^{T},\ldots,L^{{{w}}_{k}}({{z}}_{D})^{T}\right)^{T},$
and noting that the Lipschitz constant of $\bm{L}^{{{w}}_{k}}$ can be taken equal to the Lipschitz constant of $L^{{{w}}_{k}}$. To see the latter, note that for any $\bar{{z}}=\left({{z}}_{1}^{T},\ldots,{{z}}_{D}^{T}\right)^{T}$ and $\bar{{z}}^{\prime}=\left({{z}}_{1}^{\prime T},\ldots,{{z}}_{D}^{\prime T}\right)^{T}$, we have:
$\|\bm{L}^{{{w}}_{k}}(\bar{{z}})-\bm{L}^{{{w}}_{k}}(\bar{{z}}^{\prime})\|^{2}=\sum_{i=1}^{D}\|L^{{{w}}_{k}}({{z}}_{i})-L_{{{w}}_{k}}({{z}}_{i}^{\prime})\|^{2}\leq\sum_{i=1}^{D}{\mathcal{L}}_{L_{{{w}}_{k}}}^{2}\|{{z}}_{i}-{{z}}_{i}^{\prime}\|^{2}={\mathcal{L}}_{L_{{{w}}_{k}}}^{2}\|\bar{{z}}-\bar{{z}}^{\prime}\|^{2}.$
Therefore, ${\mathbb{E}}_{{{w}}_{k}\sim p_{{{w}}_{k}}}\left[{\mathcal{L}}_{\bm{L}^{{{w}}_{k}}}^{2}\right]^{1/2}\leq{\mathbb{E}}_{{{w}}_{k}\sim p_{{{w}}_{k}}}\left[{\mathcal{L}}_{L^{{{w}}_{k}}}^{2}\right]^{1/2}.$ and the same bound from Theorem 17 applies to the set-valued input case.
Proof of Theorem 21: Convergence of the ${\mathbb{W}}_{\rho}$-Error of the GMM Approximation of a SNN
We prove the theorem for the $\rho$-Wasserstein with $\rho\in 1,2$. Let us denote
$\hat{\mathbb{W}}_{2,{G}_{M_{k}}}={\mathbb{MW}}_{2}\left(q_{nn({\mathcal{X}}),k},{G}_{{M_{k}}}(q_{nn({\mathcal{X}}),k})\right),\quad\text{and}\quad\hat{\mathbb{W}}_{2,\bm{{\Delta}}_{\bm{{\mathcal{C}}}_{k}}}=\hat{\mathbb{W}}_{2,\bm{{\Delta}}_{\bm{{\mathcal{C}}}_{k}}}\left({G}_{{M_{k}}}(q_{nn({\mathcal{X}}),k})\right).$
From Corollary 20:
${\mathbb{W}}_{\rho}(p_{nn({\mathcal{X}})},q_{nn({\mathcal{X}})})$
$\qquad\leq{\mathcal{L}}_{K}\left[\hat{\mathbb{W}}_{\rho,K}+\hat{\mathbb{W}}_{2,{G}_{M_{K}}}+\hat{\mathbb{W}}_{2,\bm{{\Delta}}_{\bm{{\mathcal{C}}}_{K}}}\right]$
$\qquad={\mathcal{L}}_{K-1}{\mathcal{L}}_{K}\left[{\mathbb{W}}_{\rho,K-1}+\hat{\mathbb{W}}_{2,{G}_{M_{K-1}}}+\hat{\mathbb{W}}_{2,\bm{{\Delta}}_{\bm{}{K-1}}}\right]+{\mathcal{L}}_{K}\left[{\mathbb{W}}_{2,{G}_{M_{K}}}+{\mathbb{W}}_{2,\bm{{\Delta}}_{\bm{{\mathcal{C}}}_{K}}}\right]$
$\qquad\vdots$
$\qquad=\left(\prod_{i=1}^{K}{\mathcal{L}}_{i}\right)\left[\hat{\mathbb{W}}_{\rho,1}+\hat{\mathbb{W}}_{2,{G}_{M_{1}}}+\hat{\mathbb{W}}_{2,\bm{{\Delta}}_{\bm{{\mathcal{C}}}_{1}}}\right]+\sum_{k=2}^{K}\left(\prod_{i=k}^{K}{\mathcal{L}}_{i}\right)\left[\hat{\mathbb{W}}_{2,{G}_{M_{k}}}+\hat{\mathbb{W}}_{2,\bm{{\Delta}}_{\bm{{\mathcal{C}}}_{k}}}\right]$
$\qquad=\sum_{k=1}^{K}\left(\prod_{i=k}^{K}{\mathcal{L}}_{i}\right){\mathcal{L}}_{k}\left[\hat{\mathbb{W}}_{2,{G}_{M_{k}}}+\hat{\mathbb{W}}_{2,\bm{{\Delta}}_{{\mathcal{C}}_{k}}}\right]$
According to Corollary 15, the construction of the signature sets ${\mathcal{C}}_{k}$ in the theorem ensures that
$\hat{\mathbb{W}}_{2,\bm{{\Delta}}_{\bm{{\mathcal{C}}}_{k}}}\leq\epsilon_{k}=\frac{\epsilon}{K\prod_{i=k}^{K}{\mathcal{L}}_{i}}-\hat{\mathbb{W}}_{2,{G}_{M_{k}}}.$
Substituting this into the total bound, we get:
$\sum_{k=1}^{K}\left(\prod_{i=k}^{K}{\mathcal{L}}_{i}\right)\left[\hat{\mathbb{W}}_{2,{G}_{M_{k}}}+\hat{\mathbb{W}}_{2,\bm{{\Delta}}_{{\mathcal{C}}_{k}}}\right]$ $\leq\sum_{k=1}^{K}\left(\prod_{i=k}^{K}{\mathcal{L}}_{i}\right)\left[\hat{\mathbb{W}}_{2,{G}_{M_{k}}}+\epsilon_{k}\right]$
$=\sum_{k=1}^{K}\left(\prod_{i=k}^{K}{\mathcal{L}}_{i}\right)\left[\frac{\epsilon}{K\prod_{i=k}^{K}{\mathcal{L}}_{i}}\right]=\sum_{k=1}^{K}\frac{\epsilon}{K}=\epsilon.$
This completes the proof.
B.3 Proofs of the Results in Section 7
Proof of Corollary 26 on the ${\mathbb{W}}_{2}$ Distance between any GMM and a SNN
Since ${\mathbb{W}}_{2}$ satisfies the triangle inequality, we have:
${\mathbb{W}}_{2}\left(p_{nn({\mathcal{X}})},q_{gmm({\mathcal{X}})}\right)\leq{\mathbb{W}}_{2}\left(p_{nn({\mathcal{X}})},q_{nn({\mathcal{X}})}\right)+{\mathbb{W}}_{2}\left(q_{nn({\mathcal{X}})},q_{gmm({\mathcal{X}})}\right),$
${\mathbb{W}}_{2}\left(q_{nn({\mathcal{X}})},q_{gmm({\mathcal{X}})})\leq{\mathbb{MW}}_{2}(q_{nn({\mathcal{X}})},q_{gmm({\mathcal{X}})}\right).$
Combining these two inequalities, we obtain:
${\mathbb{W}}_{2}\left(p_{nn({\mathcal{X}})},q_{gmm({\mathcal{X}})}\right)\leq{\mathbb{W}}_{2}\left(p_{nn({\mathcal{X}})},q_{nn({\mathcal{X}})}\right)+{\mathbb{MW}}_{2}\left(q_{nn({\mathcal{X}})},q_{gmm({\mathcal{X}})}\right),$
which completes the proof.
Proof of Proposition 23 on ${\mathbb{W}}_{2}$-Optimal Grid for Signatures for Gaussians
According to Corollary 11:
${\mathbb{W}}_{2}\left({\Delta}_{\mathrm{Post}({\mathcal{C}}^{1}\times\ldots\times{\mathcal{C}}^{n},{{T}}^{-1})+\{{m}\}}\#\mathcal{N}({m},\Sigma),\mathcal{N}({m},\Sigma)\right)=\sum_{j=1}^{n}{{{\lambda}}^{(j)}}{\mathbb{W}}^{2}_{2}\left({\Delta}_{{\mathcal{C}}^{j}}\#\mathcal{N}(0,1),\mathcal{N}(0,1)\right).$